OpenStudy (anonymous):

What volume of air will be expelled from a vessel containing 500ml at 10 degreeC when it is heated to 37 degree C at the same pressure

OpenStudy (shiraz14):

Again we need to convert the temperatures to the absolute scale. 10 deg C = (10+273.15)K = 283.15K; 37 deg C = (37+273.15)K = 310.15K. Assuming ideal gas conditions, we can say that: Isobaric expansion of air = (310.15-283.15)/(283.15) x 500 ml = 47.678 ml (Answer)